Show that whether x5 + 10x3 + x + 1 is O(x4) or not?


1
Expert's answer
2021-11-15T15:57:23-0500

Let us show by contradiction that "x^5 + 10x^3 + x + 1" is not "O(x^4)". Suppose that "x^5 + 10x^3 + x + 1" is "O(x^4)". Then there exist a constant "C" and "x_0" such that "x^5 + 10x^3 + x + 1\le Cx^4" for all "x\ge x_0." It follows that "\frac{x^5 + 10x^3 + x + 1}{x^4}\le C" for all "x\ge x_0."

Since "\lim\limits_{x\to\infty}\frac{x^5 + 10x^3 + x + 1}{x^4}
=\lim\limits_{x\to\infty}(x+\frac{10}{x}+\frac{1}{x^3}+\frac{1}{x^4})=\infty," we conclude that there exists "x_1>x_0" such that "\frac{x_1^5 + 10x_1^3 + x_1 + 1}{x^4}> C." This contradiction proves that "x^5 + 10x^3 + x + 1" is not "O(x^4)".